The question is and it doesnt make sense to me.
A orange tree at 7ft tall products 18pounds of oranges. At 12ft it products 62pounds. If you minus 62-18=44, 12-7=5 so it is 44/5. Use the equation 44/5(h-7) to figure out a 18ft tree and 20ft tree and how many pounds they have.

Found 2 solutions by richwmiller, scott8148:
Answer by richwmiller(17219)

Think of the data as two points on a line.
7,18
and
12, 62


Answer by scott8148(6628)

the derived equation doesn't work for the 7ft or 12 ft tree, so why should it work for the others?

if a linear relation is correct, the 44/5 is the slope (lbs of oranges per ft of tree)
___ there also may be a "fudge factor" (y-intercept) that comes into play

18 = (44/5)(7) + ? ___ 18 = 308/5 + ? ___ -43.6 = ?

62 = (44/5)(12) - 43.6 ___ 62 = 528/5 - 43.6 ___ 62 = 105.6 - 43.6 ___ 62 = 62

(lb of oranges) = (44/5)(ft of tree) - 43.6